OpenStudy (anonymous):

I got this question wrong on my College Algebra test and i don't know why it is wrong. please help.

OpenStudy (anonymous):

Write in interval notation... \[-7 \le x < 9\] and i put \[[-7, \infty) (-\infty, 9)\] How is this wrong?

OpenStudy (anonymous):

@hartnn can you help explain this please

OpenStudy (kropot72):

Put into words, the expression says: x is equal to or greater than negative 7, and x is less than 9.

OpenStudy (jackthegreatest):

first of all, x does not go on to infinite

OpenStudy (jackthegreatest):

i just had this lesson in math

hartnn (hartnn):

which also means x is between -7 and 9 (-7 included) what you answered meant: x is greater than -7 till infinity, and x is less than 9 till -infinity

OpenStudy (kropot72):

You can find how to express this here: http://www.mathwords.com/i/interval_notation.htm

OpenStudy (jackthegreatest):

the answer should be [-7,9) i am almost certain

hartnn (hartnn):